# Searching for strong lensing by late-type galaxies in UNIONS

### v1

@@ 7 Discussion
Table 2: Summary results of the lens search. This includes all candidates regardless of how they were found. We cross-match our whole sample of lens candidates against the the Strong Lens Database (SLED) to identify the new candidates. We marginalise by grade, and present results for both strong lensing in general, and edge-on lenses only.
@@
Table 3: Summary of the performance of the CNN in identifying edge-on lenses. The metrics are marginalised by grade and magnitude bin to enable direct comparison with Table 1. We present the number of candidates found, the expected number of candidates based on Table 1, and the completeness (number found / expected number). We also include the fraction of lens candidates among sources inspected for the CNN-selected sample and for the randomly selected sample, respectively; and the ratio between these two, corresponding to the boost in lensing prevalence due to the CNN selection relative to the random selection.
@@
C 20 to 20.5 1 $<106$ $>1\%$ $2.5\times 10^{-4}$ 0
Table 4: Summary of the performance of the CNN in identifying lens candidates with any type of deflector. This table follows the layout of Table 3.
@@
Table 5: Completeness of the entire lens search, including the candidates discovered by the CNN, the candidates discovered serendipitously, and the candidates discovered during the initial characterisation of the sample. We use the estimates from Table 1 for the expected total number of candidates.
@@
We cross-match the coordinates of all the lens candidates against the Strong Lens Database (SLED, Vernardos et al., in prep.), and summarise the results in Table 2. Overall, out of the 149 candidates identified in this work (grade A, B, and C), only 17 have been reported in SLED, including two out of the 82 edge-on lens candidates. This is consistent with the idea that edge-on lenses are missed in lens searches unless they are explicitly targeted. On the other hand, given the lack of high-resolution imaging and spectroscopic data, our search is susceptible to contamination by recent mergers and random alignments of galaxies. But, the UNIONS footprint overlaps with that of Euclid, guaranteeing high-resolution imaging for all of the candidates in the next five years. This will allow us to confirm or refute many of the candidates, especially in the grade C category.
@@
Regarding our lens finding methodology, specifically the use of a CNN trained and validated on simulations, we note that the validation metrics presented in Fig. 8 overestimate the performance of the classifier when applied to the parent sample. In Fig. 8, we used the prevalence of edge-on lensing as estimated in Sect. 3, along with the recall and false positive rate of the CNN estimated on the test set, to predict 180 edge-on lenses in the 17 5141751417\,51417 514 sources selected by the network. However, only $63$ ($35\%$ of the predicted number) edge-on lenses were found during the visual inspection among the CNN-selected candidates. This implies that the use of a simulated test set leads to overestimating the purity of the network by at least a factor of two. We believe this to be a consequence of imperfect domain adaptation, and that it mainly affects $P(\text{C}\,|\,\text{L})$, the classifier’s ability to identify a real lens as a lens. Since the classifier is trained on simulations, it is expected that its performance worsens when applied to real lenses. However, this should be kept in mind when evaluating different networks and classification schemes, which is routinely done using metrics calculated on data sets dominated by mock lenses.
@@
As an alternative to the use of metrics calculated on simulated data sets, we evaluate the effectiveness of our lens search methodology by comparing the CNN-selected sample against the sample that we use in Sect. 3 to characterise the parent sample. In short, we expect the CNN-selected sample to have a higher prevalence of edge-on lenses than the parent sample, while retaining as many edge-on lenses as possible. We present these metrics in Table 3 for edge-on lenses, and in Table 4 for lensing in general. We note that using the network increases the prevalence of grade A and B edge-on lensing by up to two orders of magnitude, going from 1/30 0001300001/\$30\,000\$1 / 30 000 to $1/130$ in the brightest magnitude bin, but at the cost of recovering only $36\%$ of the lenses. However, the situation is less straightforward for the fainter edge-on lenses, since their prevalence is so low that we did not find a single instance of them when characterising the parent sample. In those cases, our estimates of the completeness are calculated against the upper bounds presented in Table 1. Nonetheless, we did find 17 edge-on lens candidates grade A and B with magnitudes between 18 and 20 in the CNN-selected sample, indicating that the CNN is sensitive to them. We also observe in the last column of Tables 3 and 4, that the improvement in the prevalence decreases with magnitude, which can be explained by the reduced contrast between the lensing features and the deflector’s light, leading to noisier classifications by the CNN. Similarly, the ratio is lower for grade C candidates, being around 100 compared with 230 for the grade A and B. This suggests that the CNN performs best against clear lens candidates, such as those with higher visual inspection grades and brighter features.
@@
Lastly, we present in Table 5 the completeness of the whole search, including all the candidates regardless of how they were identified. The search is about $64\%$ complete for the grade A and B edge-on lens candidates in brightest magnitude bin (17 to 18), and at least $19\%$ and $9\%$ complete for those brighter than 19 and 20 magnitudes, respectively. Regarding the grade C, our search finds $31\%$ of the brightest edge-on lens candidates, and around $20\%$ of those with magnitudes between 18 and 20. However, even after including the serendipitous discoveries, our search still fails to find a significant number of edge-on lens candidates in the faintest magnitude bin. Hinting that at magnitudes fainter than 20, the signal-to-noise ratio (SNR) falls below the minimum threshold at which we can identify the lensing features.
@@ 8 Conclusions
We performed a lens search targeting edge-on late-type galaxies in $3600$\,\rm{deg}^{2}$ of UNIONS $r$-band data using a CNN. We discovered 4 grade A, 20 grade B, and 58 grade C edge-on lens candidates; as well as 16 grade A, 16 grade B, and 18 grade C lens candidates that are not edge-on. This represents a significant increase in the sample of known edge-on lens candidates, previously on the order of a hundred. Furthermore, we estimate that our new candidates encompass between $9\%$ and $60\%$ of all the discoverable edge-on lens candidates in the parent sample, as determined by the initial inspection of the data in Sect. 3.
@@
We characterised the parent sample by performing a visual inspection of 120 000120000120\,000120 000 randomly selected sources. This allowed us to establish a baseline against which we can measure the performance of the CNN in the lens finding task. Moreover, using Bayes’ theorem, we extrapolated the number of candidates found among the randomly selected sources, to estimate the total number of lens candidates in the parent sample. Overall, we estimate that there are at least 8 grade A and B, and about 250 grade C edge-on lens candidates in the entire parent sample. This is consistent with a prevalence rate of one edge-on lens candidate for every 30 0003000030\,00030 000 sources. Also, for lensing in general, we find the prevalence rate to be consistent with one lens candidate for every 10 0001000010\,00010 000 sources.
@@
Regarding the performance of the CNN classifier, we showed that metrics calculated on simulated data sets significantly overestimate the performance of the network. Thus, we advise against using metrics calculated on simulations. As an alternative, we have used our initial characterisation of the sample to directly evaluate the performance of automatic classifiers without relying on simulated data sets. That is, by comparing the number of lenses and purity of the CNN-selected sample to the estimated number of lenses and prevalence of lensing in the parent sample. An ideal classifier would select all of the lenses estimated to be in the parent sample, and nothing else, resulting in a purity of 1 and a false-positive rate of 0. We propose that future lens searches follow a similar methodology when evaluating their classifiers. Reporting both the completeness of their sample and the ratio between the purity of the sample and the estimated prevalence of lensing in the parent sample.
@@
To create a training set for the CNN, we developed a data-driven methodology that allows us to systematically generate mock observations of edge-on lenses. Our mock observations use real $r$-band UNIONS observations for the deflector light, and deep HST observations for the source light, which improves the realism of the simulated images. Moreover, the simulated lenses had a realistic mass model, with a stellar-mass-to-light ratio, baryonic fraction, and total mass within the accepted ranges for edge-on lenses, ensuring the simulated lenses are physically plausible.
@@
We spectroscopically followed-up two of our edge-on lens candidates: J1121+4351 and J1559+3147. For the former, we confirmed the redshift of the deflector to be $z_{\rm d}=0.097$ and measured the redshift of the arc at $z_{\rm s}=0.246$, but given the lack of clear counter image, high-resolution imaging is still required to confirm its lensing nature. By contrast, the spectra of the lensed images of J1559+3147 showed identical features, confirming its lensing nature; we measured the redshift of the deflector at $z_{\rm d}=0.149$ and the lensed source at $z_{\rm s}=0.691$.
[…]
@@ Acknowledgements.
We thank Elodie Savary for fruitful discussions in the early stages of this work. J. A. A. B., B. C., F. C., and K. R. acknowledge support from the Swiss National Science Foundation (SNSF). Funded by a Royal Society University Research Fellowship. This project has received funding from the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme (LensEra: grant agreement No 945536). We are honoured and grateful for the opportunity of observing the Universe from Maunakea and Haleakala, which both have cultural, historical and natural significance in Hawaii. This work is based on data obtained as part of the Canada-France Imaging Survey, a CFHT large program of the National Research Council of Canada and the French Centre National de la Recherche Scientifique. Based on observations obtained with MegaPrime/MegaCam, a joint project of CFHT and CEA Saclay, at the Canada-France-Hawaii Telescope (CFHT) which is operated by the National Research Council (NRC) of Canada, the Institut National des Science de lUnivers (INSU) of the Centre National de la Recherche Scientifique (CNRS) of France, and the University of Hawaii. This research used the facilities of the Canadian Astronomy Data Centre operated by the National Research Council of Canada with the support of the Canadian Space Agency. This research is based in part on data collected at Subaru Telescope, which is operated by the National Astronomical Observatory of Japan. Pan-STARRS is a project of the Institute for Astronomy of the University of Hawaii, and is supported by the NASA SSO Near Earth Observation Program under grants 80NSSC18K0971, NNX14AM74G, NNX12AR65G, NNX13AQ47G, NNX08AR22G, 80NSSC21K1572 and by the State of Hawaii. Based on observations made with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias. Horizon 2020/2021-2025: This project has received funding from the European Unions Horizon 2020 research and innovation programme under grant agreement No 101004719 (ORP: OPTICON RadioNet Pilot). The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID #2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID #2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID #2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSFs NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honoured to be permitted to conduct astronomical research on Iolkam Duag (Kitt Peak), a mountain with particular significance to the Tohono Oodham Nation. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A&M University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de lEspai (IEEC/CSIC), the Institut de Fisica dAltes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSFs NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas A&M University. BASS is a key project of the Telescope Access Program (TAP), which has been funded by the National Astronomical Observatories of China, the Chinese Academy of Sciences (the Strategic Priority Research Program The Emergence of Cosmological Structures Grant No. XDB09000000), and the Special Fund for Astronomy from the Ministry of Finance. The BASS is also supported by the External Cooperation Program of Chinese Academy of Sciences (Grant No. 114A11KYSB20160057), and Chinese National Natural Science Foundation (Grant No. 12120101003, No. 11433005). The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract; and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
Appendix A The contrast check

### v2

@@ 7 Discussion
Table 2: Summary of the lens search results.  [^3] Includes all candidates regardless of how they were found. We cross-match our whole sample of lens candidates against the the Strong Lens Database (SLED) to identify the new candidates. We marginalise by grade, and present results for both strong lensing in general, and edge-on lenses only.
@@
Table 3: Summary of the performance of the CNN at identifying edge-on lens candidates.  [^4] Metrics are marginalised by grade and magnitude bin to enable direct comparison with Table 1. We present the number of candidates found, the expected number of candidates based on Table 1, and the completeness (number found / expected number). We also include the fraction of lens candidates among sources inspected for the CNN-selected sample and for the randomly selected sample, respectively; and the ratio between these two, corresponding to the boost in lensing prevalence due to the CNN selection relative to the random selection.
@@
C 20 to 20.5 1 <106<106 $1\%$ $2.5\times 10^{-4}$ 0
Table 4: Summary of the performance of the CNN at identifying lens candidates regardless of deflector type.  [^5] Table follows the layout of Table 3.
@@
Table 5: Overall completeness of the lens search.  [^6] Table includes candidates discovered by the CNN, candidates discovered serendipitously, and candidates discovered during the initial characterisation of the sample. We use the estimates from Table 1 for the expected total number of candidates.
@@
We cross-matched the coordinates of all the lens candidates against the Strong Lens Database (SLED, Vernardos et al., in prep.) and summarise the results in Table 2. Overall, out of the 149 candidates identified in this work (grade A, B, and C), only 17 have been reported in SLED, including two out of the 82 edge-on lens candidates. This is consistent with the idea that edge-on lenses are missed in lens searches unless they are explicitly targeted. On the other hand, given the lack of high-resolution imaging and spectroscopic data, our search is susceptible to contamination by recent mergers and random alignments of galaxies. But, the UNIONS footprint overlaps with that of Euclid, guaranteeing high-resolution imaging for all of the candidates in the next five years. This will allow us to confirm or refute many of the candidates, especially in the grade C category.
@@
Regarding our lens finding methodology, specifically the use of a CNN trained and validated on simulations, we note that the validation metrics presented in Sect. 5.2 overestimate the performance of the classifier when applied to the parent sample. In Sect. 5.2, we used the prevalence of edge-on lensing as estimated in Sect. 3, along with the recall and FPR of the CNN estimated on the test set, to predict that there would be 208 good edge-on lens candidates in the 17 51417\,514 sources selected by the network. However, only $63$ edge-on candidates ($30\%$ of the predicted number) were found during the visual inspection among the CNN-selected candidates. This implies that the use of a simulated test set leads to an overestimation of the purity of the network by at least a factor of two. We believe this is a consequence of imperfect domain adaptation, and that it mainly affects the classifier’s recall ($P(\text{C}\,|\,\text{L})$), that is, its ability to identify a real lens as a lens. Since the classifier is trained on simulations, it is expected that its performance worsens when applied to real lenses. From comparing the image properties of the mock lenses in the test set against the edge-on candidates found in this work in Appendix E, we find that the mock lenses and edge-on candidates both follow similar distributions of background noise, but the edge-on candidates have a considerable higher S/N. Moreover, the simulated mocks followed a flat Einstein radius distribution, which does not reflect the expected distribution for edge-on lenses in CFIS imaging. Overall, the observed limitations of simulated mocks should be kept in mind when evaluating different networks and classification schemes, which is routinely done using metrics calculated on datasets dominated by mock lenses.
@@
As an alternative to the use of metrics calculated on simulated datasets, we evaluate the effectiveness of our lens search methodology by comparing the CNN-selected sample against the sample that we use in Sect. 3 to characterise the parent sample. In short, we expect the CNN-selected sample to have a higher prevalence of edge-on lenses than the parent sample, while retaining as many edge-on lenses as possible. We present these metrics in Table 3 for edge-on lenses, and in Table 4 for lensing in general. We note that using the network increases the prevalence of grade A and B edge-on lensing by up to two orders of magnitude, going from 1/30 0001/\$30\,000\$ to $1/130$ in the brightest magnitude bin, but at the cost of recovering only $36\%$ of the lenses. However, the situation is less straightforward for the fainter edge-on lenses, since their prevalence is so low that we did not find a single instance of them when characterising the parent sample. In those cases, our estimates of the completeness are calculated against the upper bounds presented in Table 1. Nonetheless, we did find 17 edge-on lens candidates grade A and B with magnitudes between 18 and 20 in the CNN-selected sample, indicating that the CNN is sensitive to them. We also observe in the last column of Tables 3 and 4, that the improvement in the prevalence decreases with magnitude, which can be explained by the reduced contrast between the lensing features and the deflector’s light, leading to noisier classifications by the CNN. Similarly, the ratio is lower for grade C candidates, being around 100 compared with 230 for the grade A and B. This suggests that the CNN performs best against clear lens candidates, such as those with higher visual inspection grades and brighter features.
@@
Moreover, we present in Table 5 the completeness of the whole search, including all the candidates regardless of how they were identified. The search is about $64\%$ complete for the grade A and B edge-on lens candidates in brightest magnitude bin (17 to 18), and at least $19\%$ and $9\%$ complete for those brighter than 19 and 20 magnitudes, respectively. Regarding the grade C, our search finds $31\%$ of the brightest edge-on lens candidates, and around $20\%$ of those with magnitudes between 18 and 20. However, even after including the serendipitous discoveries, our search still fails to find a significant number of edge-on lens candidates in the faintest magnitude bin. Hinting that at magnitudes fainter than 20, the S/N falls below the minimum threshold at which we can identify the lensing features.
Finally, having observed the limitations of metrics calculated on simulations, and the fact that the empirical metrics proposed in this work are only available after running the lens search, we propose an iterative approach to improve the purity of lens searches. In this approach, networks are first trained on simulated mocks plus the few lenses already known in the parent sample. Then, after visually inspecting the network-selected sample and curating a final list of candidates, the networks should be retrained using the full sample of lens candidates available. Given that the retraining sample will be on the order of tens of, or at most a few hundred candidates, the retraining must be done carefully, either by freezing most layers and retraining only the final ones or by combining the new lens candidates with the original simulations and gradually removing the simulations from the training data epoch by epoch, leaving only the lens candidates by the end. Once the model has been retrained, the empirical metrics should be recalculated and used to assess its performance. Since the prevalence of lensing has already been estimated to evaluate the first model, the calculation of the empirical metrics will be more efficient and less time-consuming for the retrained models.
@@ 8 Conclusions
We have performed a lens search targeting edge-on late-type galaxies in $3600$\,\rm{deg}^{2}$ of UNIONS $r$-band data using a CNN. We discovered 4 grade A, 20 grade B, and 58 grade C edge-on lens candidates as well as 16 grade A, 16 grade B, and 18 grade C lens candidates that are not edge-on. This represents a significant increase in the sample of known edge-on lens candidates, which was previously on the order of one hundred. Furthermore, we estimate that our new candidates encompass between $9\%$ and $60\%$ of all the discoverable edge-on lens candidates in the parent sample, as determined by the initial inspection of the data in Sect. 3.
@@
We characterised the parent sample by performing a visual inspection of 120 000120\,000 randomly selected sources. This allowed us to establish a baseline against which we can measure the performance of the CNN in the lens finding task. Moreover, using Bayes’ theorem, we extrapolated the number of candidates found among the randomly selected sources to estimate the total number of lens candidates in the parent sample. Overall, we estimate that there are at least eight grade A and B edge-on lens candidates and about 250 grade C edge-on lens candidates in the entire parent sample. This is consistent with a prevalence rate of one edge-on lens candidate for every 30 00030\,000 sources. Also, for lensing in general, we find the prevalence rate to be consistent with one lens candidate for every 10 00010\,000 sources.
@@
Regarding the performance of the CNN classifier, we show that metrics calculated on simulated datasets significantly overestimate the performance of the network. Thus, we advise against using metrics calculated on simulations. As an alternative, we used our initial characterisation of the sample to directly evaluate the performance of our classifier without relying on simulated datasets. That is, we compared the number of lenses and the purity of the CNN-selected sample to the estimated number of lenses and the prevalence of lensing in the parent sample. An ideal classifier would select all of the lenses estimated to be in the parent sample and nothing else, resulting in a purity of one and a false-positive rate of zero. We propose that future lens searches follow a similar methodology when evaluating their classifiers, reporting both the completeness of their sample and the ratio between the purity of the sample and the estimated prevalence of lensing in the parent sample. Moreover, we also propose the optimisation of these metrics when retraining models on newly discovered lens candidates.
@@
To create a training set for the CNN, we developed a data-driven methodology that allows us to systematically generate mock observations of edge-on lenses. Our mock observations use real $r$-band UNIONS observations for the deflector light and deep HST observations for the source light, which improves the realism of the simulated images. Moreover, the simulated lenses had a realistic mass model, with a stellar-mass-to-light ratio, baryonic fraction, and total mass within the accepted ranges for edge-on lenses, ensuring the simulated lenses are physically plausible.
@@
We spectroscopically followed up on two of our edge-on lens candidates: J1121+4351 and J1559+3147. For the former, we confirmed the redshift of the deflector to be $z_{\rm d}=0.097$ and measured the redshift of the arc at $z_{\rm s}=0.246$, but given the lack of a clear counter image, high-resolution imaging is still required to confirm its lensing nature. By contrast, the spectra of the lensed images of J1559+3147 showed identical features, confirming its lensing nature; we measured the redshift of the deflector at $z_{\rm d}=0.149$ and the lensed source at $z_{\rm s}=0.691$.
Finally, we prove the feasibility of finding edge-on lenses in large-area optical surveys using artificial intelligence, as long as the lens search targets edge-on lenses explicitly. However, even though the UNIONS $r$-band data have an excellent average seeing (0.6), we hit the limitations of ground-based observations. The lower lensing cross-section of late-type galaxies, combined with their extended spatial features, results in most lens candidates suffering from blending between the light of the deflector and the lensing features. This makes it challenging for human experts to identify lens candidates during visual inspection and can also confuse CNN classifiers. This is evident in the decrease in the number of candidates found as the magnitude increases. Thus, we conclude that to identify the bulk of the late-type galaxy lenses, it is necessary to mine surveys with an even higher resolution. This is already the case with HSC, which was providing most of the new edge-on late-type lens candidates before this work. However, the most significant step forward will come from Euclid, which will observe $14\,000$\,\deg^{2}$ with an average resolution of 0.2, combining a large volume probed with superb resolution, thus making it the most promising target for future searches of lensing by edge-on late-type galaxies.
Appendix A Acknowledgements
@@ Acknowledgements.
We thank Elodie Savary for fruitful discussions in the early stages of this work. J. A. A. B., B. C., F. C., and K. R. acknowledge support from the Swiss National Science Foundation (SNSF). Funded by a Royal Society University Research Fellowship. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LensEra: grant agreement No 945536). We are honoured and grateful for the opportunity of observing the Universe from Maunakea and Haleakala, which both have cultural, historical and natural significance in Hawaii. This work is based on data obtained as part of the Canada-France Imaging Survey, a CFHT large program of the National Research Council of Canada and the French Centre National de la Recherche Scientifique. Based on observations obtained with MegaPrime/MegaCam, a joint project of CFHT and CEA Saclay, at the Canada-France-Hawaii Telescope (CFHT) which is operated by the National Research Council (NRC) of Canada, the Institut National des Science de l’Univers (INSU) of the Centre National de la Recherche Scientifique (CNRS) of France, and the University of Hawaii. This research used the facilities of the Canadian Astronomy Data Centre operated by the National Research Council of Canada with the support of the Canadian Space Agency. This research is based in part on data collected at Subaru Telescope, which is operated by the National Astronomical Observatory of Japan. Pan-STARRS is a project of the Institute for Astronomy of the University of Hawaii, and is supported by the NASA SSO Near Earth Observation Program under grants 80NSSC18K0971, NNX14AM74G, NNX12AR65G, NNX13AQ47G, NNX08AR22G, 80NSSC21K1572 and by the State of Hawaii. Based on observations made with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias. Horizon 2020/2021-2025: This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101004719 (ORP: OPTICON RadioNet Pilot). The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID #2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID #2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID #2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honoured to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A&M University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de l’Espai (IEEC/CSIC), the Institut de Fisica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas A&M University. BASS is a key project of the Telescope Access Program (TAP), which has been funded by the National Astronomical Observatories of China, the Chinese Academy of Sciences (the Strategic Priority Research Program “The Emergence of Cosmological Structures” Grant No. XDB09000000), and the Special Fund for Astronomy from the Ministry of Finance. The BASS is also supported by the External Cooperation Program of Chinese Academy of Sciences (Grant No. 114A11KYSB20160057), and Chinese National Natural Science Foundation (Grant No. 12120101003, No. 11433005). The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract; and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
Appendix B Lens candidates as seen by UNIONS
Figure 7: Edge-on lens candidates identified by the trained CNN or discovered serendipitously either by prototypes of the classifier or while preparing the mock lenses. They correspond to 5 grade A, 19 grade B, and 50 grade C edge-on lens candidates, and of these, only two are already known. The layout follows the same scheme from Fig. 2, with the addition of a rectangle outlined with a red dotted line highlighting the position of the slit for the lens candidates with spectroscopic follow-up.
Figure 8: Non-edge-on lens candidates identified by the trained CNN or discovered serendipitously either by prototypes of the classifier or while preparing the mock lenses. The layout follows the same scheme from Fig. 2. The 38 candidates correspond to 22 grade A, 12 grade B, and 4 grade C. Of these, 12 are reported in SLED.
Appendix C The contrast check